Kackar Mountain gets 1.8x103 km3 snow at wintertime. Once spring comes and all the snow gets melt, it would result from the additional luqiuid water entering the Black Sea. The are of Black Sea is around 4.36x105 km2 and the density of ice and water are 0.92 g/cm3 and 1.0 g/cm3, respectively. Estimate the rise in sea level, if all the snow melts.


1
Expert's answer
2021-10-24T01:32:49-0400

1.8cm*/0.92=1.656*10^3=1.656cm

4.36*10^2/1=4.36*10^5=436cm

1.656cm+436cm=437,656cm